OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 7, -7, 7, -7, ...

OpenStudy (anonymous):

I don't think there's a rule since it's 7, -7, and it keeps going like that forever

OpenStudy (anonymous):

It diverges

OpenStudy (anonymous):

I had a problem like this before and it was 1, -1, 1, -1... and I wrote that it diverges and got it right

OpenStudy (anonymous):

@jennisicle its not in my answer choices an = 7 • (-1)n - 1 an = 7 • (-1)n an = 7 • 1n - 1 an = 7 • 1n + 1

OpenStudy (anonymous):

a1 = 7, a2 = -7, etc an = 7(-1)^(n-1) n = 1 to ... A